OpenStudy (anonymous):

im having ahard time with this question by trial and error with your calculator, find the angle theta for which sin theta and theta differ by 5%. This calculation must be done in radians. when you have found theta, express it in degrees.

OpenStudy (anonymous):

First, let's define percent difference. \[{\rm Percent~Difference} = {\rm difference \over maximum}\]where difference is the difference between \(\sin \theta\) and \(\theta\) and maximum is the value that has the greatest magnitude. Take a look here for more information on small-angle approximations. http://en.wikipedia.org/wiki/Small-angle_approximation \[0.05 = {|\sin(\theta) - \theta| \over \theta}\]Since \(\theta\) will be greater than \(\sin(\theta)\) for all values of \(\theta \gt 0\) There is no direct method to solve this. You're going to have to do a little guess and check. Might I recommend starting somewhere less than \(\theta \lt 1\)

OpenStudy (anonymous):

umm can u tell me if 0.02 is right?

OpenStudy (anonymous):

when i put in the radians i got 0.019

OpenStudy (anonymous):

I'm getting something close to 0.5 radians.

OpenStudy (anonymous):

the way im doing is i set up it like this sintheta= 0.95(theta)

OpenStudy (anonymous):

and my calculator is is in radians

OpenStudy (anonymous):

Be careful with that expression. That assumes that theta must be 5% less than sin(theta). However, percent difference says that sin(theta) can be 5% less than theta. Since we know that sin(theta) < theta for all theta, lets use this expression. \[0.95 \sin \theta = \theta\]

OpenStudy (anonymous):

sin (0.5) =0.479 radians 0.5/0.95=0.526

OpenStudy (anonymous):

That looks like what I got.

OpenStudy (anonymous):

Within 1%

OpenStudy (anonymous):

so to get the degree do i just say 0.5 radians (180 degrees)/radians

OpenStudy (anonymous):

\[\rm degrees = {180 \over \pi}\]

OpenStudy (anonymous):

so its not 0.5 radians time 180 degrees/ radians

OpenStudy (anonymous):

is the angle 90

OpenStudy (anonymous):

Negative. The conversion is\[\rm degrees = {180 \over \pi}\] You should get approximately 28.5.

OpenStudy (anonymous):

how canu please show me

OpenStudy (anonymous):

Sorry. \[\rm degress = [radians] {180 \over \pi}\]

OpenStudy (anonymous):

so 0.5 is not treated as radians because i thought it was 0.5pi(180/pi)

OpenStudy (anonymous):

the question asks the difference betwene theta and sin theta should only be 5 percent but in my situation ios only 1

OpenStudy (anonymous):

I think I'm confused as to where you are confused. We solved for theta such that a 5% difference had to be satisfied.

OpenStudy (anonymous):

okay just to make sure i calculated properly i did o.526-0.479=0.47x100 =5percent

OpenStudy (anonymous):

so the difference is 5 percent

OpenStudy (anonymous):

i read the lin u gave me but im still not sure how u knew that theta had to be leass than one (srry eashmore for bithering u )

OpenStudy (anonymous):

I know because I've used this approximation several times. Also note that this approximation is called the "small-angle approximation." I also ran the numbers, so I know where the answer should be and I wanted to guide you in the right direction. I might suggest doing one more iteration of your calculations. Try 0.55 radians. I think you'll be pleasantly surprised.

OpenStudy (anonymous):

difference is ab0ut 5.6 percent.. would it be wrong if we just looked at theta vs. sin theta. rather than sin theta vs. theta /5

OpenStudy (mani_jha):

http://openstudy.com/users/anna818#/updates/4f68107de4b0f81dfbb53ce7 I did it that way.....